OpenStudy (anonymous):

what is the net force and acceleration of a ball of mass 0.750 kilograms that is subjected to forces 11.00 newtons and 7.00 newtons from opposite directions?

OpenStudy (anonymous):

F=m*a. We know m, and are solving for a. F1=11 Newtons and F2=-7 Newtons (negative because they're opposite in direction). We can sum the forces up, F1+F2= 4 Newtons = .750 kg * a. now solve for acceleration, a=4/.750 and that should be the acceleration. The net force is just the F1+F2=4 Newtons.